The probability of getting heads on a biased coin is 0.8. You flip the coin twice. What is the probability of getting one each of heads and tails?

Step 1: Calculate probability of getting tails by subtracting 1-0.8 = 0.2.Step 2: Calculate probability of getting heads then tails by multiplying 0.80.2 = 0.16.Step 3: Calculate probability of getting tails then heads by multiplying 0.20.8 = 0.16. Step 4: Calculate total probability of getting one each of heads and tails by adding 0.16+0.16=0.32 // OR recognise that P(heads then tails)=P(tails then heads) and calculate total probability by multiplying 0.16*2 = 0.32.You can also use a tree diagram to help illustrate the different possible outcomes.

NR

Related Maths GCSE answers

All answers ▸

Solve the following set of equations. 3x + 2y = 5, 2x + 3y =6


How do I factorise x^2 + 8x + 15?


Calculate 7.6 x 32.5


Tommy, Anna and Jacob all have 40 sweets. they decide to split the sweets between each other in the ratio 1:4:5. Calculate how many sweets each get, rounding down your answer where necessary.